OpenStudy (anonymous):

medals rewarded! Evaluate the function y=-3(1/4)^x. solve for x=-3

OpenStudy (anonymous):

your job is to compute \[3\left(\frac{1}{4}\right)^3\]

OpenStudy (anonymous):

you should get \(\frac{3}{64}\)

OpenStudy (jhannybean):

-3rd power...

OpenStudy (anonymous):

oh, does it say \(-3\) ??

OpenStudy (jhannybean):

Same thing.(1/4)^{-3} = 4^3.

OpenStudy (anonymous):

the take the reciprocal and get \[3\times 64\]

OpenStudy (jhannybean):

3(1/4)^{-3) = 3(4^3)

OpenStudy (anonymous):

I got 192, is that correct

OpenStudy (jhannybean):

Yep, good job.